# JointMotion: Joint Self-supervision for Joint Motion Prediction

### v1

@@ 4.2 Comparison of self-supervised pre-training methods for joint motion prediction
In this experiment, we compare our method with recent self-supervised pre-training methods for autonomous driving using the WOMD dataset. Specifically, we compare with contrastive learning via PreTraM [33] and two masked autoencoding approaches, Forecast-MAE [9] and Traj-MAE [7].
@@
Experimental setup. We pre-train and fine-tune well-established Scene Transformer [22] models on the WOMD training split. We use the publicly available implementation by [36] with 3 modality-specific encoders (i.e., for agents, lanes, and traffic lights) and a shared decoder for $K=6$ motion modes. For pre-training, we add an shared local decoder for late fusion models (see Fig. 2) with 3 transformer blocks and the hyperparamters described in Sec. 3.2. For our method, we additionally add two projectors for scene-level representations as described in Sec. 3.1. For PreTraM, we follow its trajectory-map contrastive learning configuration and add two linear projection layers as projectors for trajectory and map embeddings. For Forecast-MAE, we use a masking ratio of 60% and reconstruct positions of lane polylines and past trajectories by minimizing the MSE loss. We exclude future trajectories, since self-supervised learning by design does not use the same labels as the intended downstream task (cf., [1]). For Traj-MAE, we use a masking ration of 60% and reconstruct positions of lane polylines and past trajectories by minimizing the corresponding Huber loss. For our method, we minimize the joint loss of our proposed objectives: $\mathcal{L}_{\text{JointMotion}}=\lambda_{\text{con}}\mathcal{L}_{\text{con}}+\mathcal{L}_{\text{recon}}$. We set $\lambda_{\text{con}}=0.01$ and following [35] the weight of the redundancy reduction term $\lambda_{\text{red}}=0.005$. We fine-tune the model using its joint configuration and hard loss assignment. Accordingly, the loss is computed for the best scene-wide joint prediction mode. As post-processing, we follow [19] and adjust the predicted confidences of redundant predictions. For all methods, we perform pre-training for 10 hours and fine-tuning for 23.5 hours using a training server with 4 A100 GPUs. For pre-training and fine-tuning, we use AdamW with an initial learning rate of 1e-4 and a step learning rate scheduler with a reduction rate of 0.5 and a step size of 25 epochs.
@@
Results. Tab. 1 shows the results of this experiments. Explicit scene-level objectives (i.e., PreTraM and JointMotion) lead to better and more balanced performance across all agent types, while implicit scene-wide masked autoencoding with Forecast-MAE or Traj-MAE tends to focus more on the pedestrian class than on the others (see mAP scores). Therefore, it is likely that explict scene-wide objectives enforce learning interactions between varying agent types more. Overall, our method achieves the best scores across all agent types. Our method without the objective of connecting motion and environments (JointMotion w/o CME), achieves worse performance. This shows that the combination of both objective works best and that the objectives are complementary. Fig. 3 shows the validation reconstruction losses of both configurations during pre-training and further highlights the complementary nature of our two objectives. Specifically, the reconstruction of past traffic light sequences and past motion are learned similar with both configurations. However, the lane reconstructions are more accurate without our scene-level objective. We hypothesize that models pre-trained with our scene-level objective tend to focus more on the overall lane structure than on specific details of individual lane polylines.
Table 1: Comparison of self-supervised pre-training methods for joint motion prediction. All methods are used to pre-train Scene Transformer models [22] on the Waymo Open Motion dataset and evaluated on the validation split. Agent types: cyclist (cyc), pedestrian (ped), and vehicle (veh). “w/o CME” denotes our method without the objective of connecting motion and environments. Best scores are bold, second best are underlined.
@@
None 0.1596 0.1465 0.1821 0.1504 1.522 0.6982 1.486 3.653 1.654 3.476
Forecast-MAE [9] 0.1592 0.1420 0.1873 0.1482 1.529 0.6940 1.423 3.664 1.652 3.343
Traj-MAE [7] 0.1677 0.1492 0.1917 0.1623 1.421 0.7084 1.338 3.320 1.647 3.090
PreTraM [33] 0.1689 0.1724 0.1775 0.1569 1.488 0.7108 1.461 3.489 1.657 3.374
JointMotion w/o CME 0.1784 0.1652 0.1903 0.1796 1.457 0.6996 1.317 3.363 1.630 3.033
JointMotion 2 Tr 0.3 w 0.1940  0 Tr 0 w 2 Tr 0.3 w 0.1970  0 Tr 0 w 2 Tr 0.3 w 0.1964  0 Tr 0 w 2 Tr 0.3 w 0.1886  0 Tr 0 w 2 Tr 0.3 w 1.343  0 Tr 0 w 2 Tr 0.3 w 0.6767  0 Tr 0 w 2 Tr 0.3 w 1.288  0 Tr 0 w 2 Tr 0.3 w 3.095  0 Tr 0 w 2 Tr 0.3 w 1.583  0 Tr 0 w 2 Tr 0.3 w 2.941  0 Tr 0 w
@@
Figure 3: Loss plots of complementary pre-training objectives. The green curve represents JointMotion w/o CME, while the blue curve represents JointMotion with the objective of connecting motion and environments. Consistent with the remainder of the document, L stands for lanes, TL stands for traffic lights, and A stands for agents.
@@
Figure 4: Accelerating and improving training via SSL. Scene Transformer models pre-trained with JointMotion achieve higher mAP scores on WOMD than models trained from scratch, even in the same total wall training time (pre-training + fine-tuning).
@@
Fig. 4 shows that Scene Transformer models pre-trained with our method achieve higher mAP scores on WOMD than models trained from scratch, even in the same total wall training time (pre-training + fine-tuning).
5 Comparing scene-level self-supervision methods
@@
In this experiment, we further compare the scene-level objectives PreTraM [33] and JointMotion. We train Scene Transformer, HPTR, and a joint configuration of Wayformer to cover all common types of environment representations in motion prediction (i.e., agent-centric, scene-centric, and pairwise relative).
@@
Experimental setup. For Scene Transformer, we use the same configuration as in the previous experiment. For HPTR, we add 3 modality-specific encoders and use a shared decoder for $K=6$ motion modes. For the joint configuration of Wayformer, we follow [18] and use a shared encoder for early fusion, which compresses the multi-modal input to 128 tokens. We concatenate multiple such agent-centric embeddings with positional and rotation information into a common reference frame and use a shared decoder to predict scene-wide joint modes. For pre-training the Wayformer model, we add our decoder for early fusion configurations (see Fig. 2). For all models, we employ the same optimizer and learning rate scheduling as in the previous experiment. All models are pre-trained for 10 hours and fine-tuned for 23.5 hours using a training server with 4 A100 GPUs. Evaluation is performed on the interactive validation split of WOMD and AV2. Our comparison is structured according to the performance metrics mAP (mean Average Precision), minSADE (minimum Social Average Displacement Error), minSFDE (minimum Social Final Displacement Error), MR (Miss Rate), and OR (Overlap Rate). minSADE measures the displacement error between the ground-truth future trajectories of interacting agents and the closest joint prediction out of a predefined set, averaged over the future time horizon [6]. This metric emphasizes the model’s ability to account for social interactions in predicting the average trajectory displacement accurately. On the other hand, minSFDE focuses on the displacement error at the time horizon endpoint.
@@
Results. Table 2 shows the results of this experiment. Our method consistently outperforms PreTraM using different models with varying environment representations. Our method improves all models, while the improvement of the HPTR model (e.g., 11% lower minSFDE) is most significant and the improvement of the agent-centric Wayformer is least significant (e.g., 3% lower minSFDE). The Wayformer model is not pre-trainable with PreTraM since an instance-level objective is required to decode the modality-specific tokens from fused representations (cf., Sec. 3.2). The lowest block shows that our method leads to transferable representations. Specifically, pre-training on WOMD improves fine-tuning on AV2.
@@
Table 2: Comparing scene-level self-supervision methods. All metrics are computed using the Waymo Open Motion interactive (WOMD) and Argoverse 2 Forecasting (AV2) validation splits. Best scores are bold, second best are underlined.
Dataset Pre-training Model mAP $\uparrow$ minSADE $\downarrow$ minSFDE $\downarrow$ MR $\downarrow$ OR $\downarrow$
WOMD None Scene Transformer 0.0822 1.5255 3.6715 0.7372 0.2868
PreTraM [33] Scene Transformer 0.0722 1.5415 3.6508 0.7385 0.2915
JointMotion Scene Transformer 2 Tr 0.3 w 0.0876  0 Tr 0 w 2 Tr 0.3 w 1.3830  0 Tr 0 w 2 Tr 0.3 w 3.2400  0 Tr 0 w 2 Tr 0.3 w 0.7090  0 Tr 0 w 2 Tr 0.3 w 0.2847  0 Tr 0 w
WOMD None HPTR 0.1033 1.1682 2.6003 0.6030 0.2331
PreTraM [33] HPTR 0.1286 1.0981 2.5049 0.5863 0.2345
JointMotion HPTR 2 Tr 0.3 w 0.1456  0 Tr 0 w 2 Tr 0.3 w 1.0564  0 Tr 0 w 2 Tr 0.3 w 2.4006  0 Tr 0 w 2 Tr 0.3 w 0.5591  0 Tr 0 w 2 Tr 0.3 w 0.2297  0 Tr 0 w
WOMD None Wayformer (joint) 0.1533 1.0209 2.3529 0.5461 0.2273
JointMotion Wayformer (joint) 2 Tr 0.3 w 0.1553  0 Tr 0 w 2 Tr 0.3 w 0.9939  0 Tr 0 w 2 Tr 0.3 w 2.2823  0 Tr 0 w 2 Tr 0.3 w 0.5270  0 Tr 0 w 2 Tr 0.3 w 0.2143  0 Tr 0 w
AV2 None HPTR 0.2554 2.255 1.138 - 2 Tr 0.3 w 0.0988  0 Tr 0 w
JointMotion WOMD HPTR 2 Tr 0.3 w 0.2697  0 Tr 0 w 2 Tr 0.3 w 2.153  0 Tr 0 w 2 Tr 0.3 w 1.137  0 Tr 0 w - 0.1025
6 Comparison with state-of-the-art methods for joint motion prediction
@@
Experimental setup. We pre-train HPTR (configured as in Sec. 5) for 10 hours using JointMotion and fine-tune for 100 hours. We evaluate the methods in test and validation splits of WOMD. We use the official challenge website to compute performance metrics.
Results. Table 3 presents a comparative analysis of various state-of-the-art methods for joint motion prediction on interactive splits of WOMD. In the test split, we compare four different approaches. The Scene Transformer has a mAP of 0.1192 and is outperformed by the other methods in all metrics. GameFormer achieves a higher mAP of 0.1376 and exhibits competitive performance in terms of minSADE and minSFDE. MotionDiffuser and JointMotion (HPTR) are close contenders, with MotionDiffuser showing slightly better in all of the metrics compared to JointMotion (HPTR).
For the Validation split, the results are similar. GameFormer (joint) scores a mAP of 0.1339, with MotionLM (single replica) slightly outperforming it in terms of mAP, achieving 0.1687. JointMotion (HPTR) presents an improved mAP of 0.1761 compared to its counterparts. Following HPTR, we compare against single replica model of MotionLM and show its ensembling version for reference. MotionLM (ensemble) demonstrates superior overall performance, a result anticipated given power of ensembling.
Table 3: Comparison with state-of-the-art methods for joint motion prediction. All methods are evaluated on interactive splits of the Waymo Open Motion dataset. Following [36] we compare against single replica versions of joint prediction methods, ensemble versions are shown for reference. Best scores are bold, second best are underlined.
Split Method (config) Venue mAP $\uparrow$ minSADE $\downarrow$ minSFDE $\downarrow$ MR $\downarrow$ OR $\downarrow$
@@
GameFormer (joint) [17] ICCV’23 0.1376 0.9161 2 Tr 0.3 w 1.9373  0 Tr 0 w 0.4531 0.2112
MotionDiffuser [18] CVPR’23 2 Tr 0.3 w 0.1952  0 Tr 0 w 2 Tr 0.3 w 0.8642  0 Tr 0 w 1.9482 2 Tr 0.3 w 0.4300  0 Tr 0 w 2 Tr 0.3 w 0.2004  0 Tr 0 w
@@
Val GameFormer (joint) [17] ICCV’23 0.1339 2 Tr 0.3 w 0.9133  0 Tr 0 w 2 Tr 0.3 w 1.9251  0 Tr 0 w 2 Tr 0.3 w 0.4564  0 Tr 0 w -

### v2

@@ 4.1 Comparison of self-supervised pre-training methods for joint motion prediction
In this experiment, we compare our method with recent self-supervised pre-training methods for autonomous driving using the WOMD dataset. Specifically, we compare with contrastive learning via PreTraM [15] and masked autoencoding with Forecast-MAE [17] and Traj-MAE [16].
Motion prediction models. We pre-train and fine-tune well-established Scene Transformer [20] models on the WOMD training split. We use the publicly available implementation by [21] with 3 modality-specific encoders (i.e., for agents, lanes, and traffic lights). Adapted to the complexity per modality, we encode traffic light features with 1, agent features with 3, and lane features with 6 transformer blocks.
@@
Pre-training. For all methods, we add a pre-training decoder for late fusion models (see Figure 2) with 3 transformer blocks and the hyperparameters described in Section 3.2. For our method, we additionally add two projectors for scene-level representations as described in Section 3.1. For PreTraM, we follow its trajectory-map contrastive learning configuration and add two linear projection layers as projectors for trajectory and map embeddings. For Forecast-MAE, we use a masking ratio of 60% and reconstruct positions of lane polylines and past trajectories by minimizing the MSE loss. We exclude future trajectories, since self-supervised learning by design does not use the same labels as the intended downstream task (cf. [1]). For Traj-MAE, we use a masking ratio of 60% and reconstruct positions of lane polylines and past trajectories by minimizing the corresponding Huber loss. For our method, we minimize the joint loss of our proposed objectives $\mathcal{L}_{\text{JointMotion}}=\lambda_{\text{CME}}\mathcal{L}_{\text{CME}}+\mathcal{L}_{\text{MPM}}$. We set $\lambda_{\text{CME}}=0.01$ and following [31] the weight of the redundancy reduction term $\lambda_{\text{red}}=0.005$.
Fine-tuning. For all methods, we replace the pre-training decoder with a shared global decoder and learned anchors for $k=6$ motion modes. We initialize the modality-specific encoders with the learned weights from pre-training and do not freeze any weights during fine-tuning. We fine-tune the model using its joint configuration and hard loss assignment. Accordingly, the loss is computed for the best scene-wide joint prediction mode. As post-processing, we follow [36] and adjust the confidences of redundant predictions.
Training time, hardware, and optimizer. For all methods, we perform pre-training for 10 hours and fine-tuning for 23.5 hours using a training server with 4 A100 GPUs. For pre-training and fine-tuning, we use AdamW [37] with an initial learning rate of 1e-4 and a step learning rate scheduler with a reduction rate of 0.5 and a step size of 25 epochs.
@@
Results. Table 1 shows the results of this experiments. Explicit scene-level objectives (i.e., PreTraM and JointMotion) lead to better and more balanced performance across all agent types, while implicit scene-wide masked autoencoding with Forecast-MAE or Traj-MAE tends to focus more on the pedestrian class than on the others (see mAP scores). Therefore, it is likely that explict scene-wide objectives enforce learning interactions between varying agent types more. Traj-MAE and our method without the objective of connecting motion and environments (JointMotion w/o CME) achieve better scores than Forecast-MAE. Consequently, reconstructing individual elements of polylines improves representations more than reconstructing whole polylines. The superior performance of JointMotion w/o CME compared to Traj-MAE indicates that extending the motion feature set and reconstructing traffic light states as well further improves learned representations. Our method without masked polyline modeling (JointMotion w/o MPM) performs on par with PreTraM, indicating that redundancy reduction can replace negative examples for scene-level similarity learning.
@@
None 0.1596 0.1465 0.1821 0.1504 1.522 0.698 1.486 3.653 1.654 3.476
Forecast-MAE [17] 0.1592 0.1420 0.1873 0.1482 1.529 0.694 1.423 3.664 1.652 3.343
Traj-MAE [16] 0.1677 0.1492 0.1917 0.1623 1.421 0.708 1.338 3.320 1.647 3.090
PreTraM [15] 0.1689 0.1724 0.1775 0.1569 1.488 0.711 1.461 3.489 1.657 3.374
JointMotion w/o MPM 0.1689 0.1762 0.1777 0.1528 1.478 0.684 1.406 3.507 1.608 3.287
JointMotion w/o CME 0.1784 0.1652 0.1903 0.1796 1.457 0.700 1.317 3.363 1.630 3.033
JointMotion 0.1940 0.1970 0.1964 0.1886 1.343 0.677 1.288 3.095 1.583 2.941
Table 1: Comparison of self-supervised pre-training methods for joint motion prediction. All methods are used to pre-train Scene Transformer models [20] on the Waymo Open Motion dataset and evaluated on the validation split. Agent types: cyclist (cyc), pedestrian (ped), and vehicle (veh). Best scores are bold, second best are underlined.
Overall, pre-training with both proposed objectives (i.e., JointMotion) leads to the best scores across all agent types. This shows that the combination of our objectives works best. Figure 3 further highlights the complementary nature of our two objectives. Specifically, the reconstruction of past traffic light sequences are learned very similar with both configurations. The reconstruction loss for past agent motion converges more slowly with JointMotion pre-training, but reaches similar values as well. However, with our scene-level objective, the lane reconstruction loss likely converges to a higher value. We hypothesize that models pre-trained with our scene-level objective tend to focus more on the overall lane structure than on specific details of individual lane polylines.
@@
Figure 3: Loss plots of our complementary pre-training objectives. The green curve represents JointMotion w/o CME, while the blue curve represents JointMotion. Consistent with the remainder of the document, L stands for lanes, TL stands for traffic lights, and A stands for agents.
@@
Figure 4: Accelerating and improving training via SSL. Scene Transformer models pre-trained with JointMotion achieve higher mAP scores on WOMD than models trained from scratch.
@@
Figure 4 shows that Scene Transformer models pre-trained with our method achieve higher mAP scores on WOMD than models trained from scratch, even in a shorter wall training time.
4.2 Comparing scene-level self-supervision methods
@@
In this experiment, we further compare the scene-level objectives PreTraM [15] and JointMotion. We train Scene Transformer, HPTR, and a joint configuration of Wayformer to cover all common types of environment representations in motion prediction (i.e., scene-centric, pairwise relative, and agent-centric).
@@
Experimental setup. For Scene Transformer, we use the same configuration as in the previous experiment. For HPTR, we analogously add 3 modality-specific encoders and use a shared decoder for $k=6$ motion modes. For the joint configuration of Wayformer, we follow [24] and use a shared encoder for early fusion, which compresses the multi-modal input to 128 tokens. We concatenate multiple such agent-centric embeddings with positional and rotation information into a common reference frame and use a shared decoder to predict joint motion modes. For pre-training the Wayformer model, we add our decoder for early fusion configurations (see Figure 2). The Wayformer model is not pre-trainable with PreTraM since an instance-level objective is required to decode the modality-specific tokens from fused representations (cf. Section 3.2). For all models, we employ the same hardware, training time, optimizer, and learning rate scheduling as in the previous experiment. Evaluation is performed on the interactive validation split of WOMD and AV2.
@@
Results. Table 2 shows the results of this experiment. Our method consistently outperforms PreTraM using different models with varying environment representations and fusion mechanisms. Our method improves all models, while the improvement of the scene-centric Scene Transformer model (e.g., 12% lower minFDE) is most significant and the improvement of the agent-centric Wayformer model is least significant (e.g., 3% lower minFDE). Hence, the improvements are inversely proportional to the sample efficiency of the models. In scene-centric modeling, one sample is generated per scene, whereas in agent-centric modeling, one sample is generated for each traffic agent within a scene. This aligns with the finding that fine-tuning with more samples generally reduces the value of pre-training [38].
Dataset Model (config) Pre-training minFDE $\downarrow$ minADE $\downarrow$ MR $\downarrow$ OR $\downarrow$
WOMD Scene Transformer None 3.6715 1.5255 0.7372 0.2868
PreTraM [15] 3.6508 -0.57% 1.5415 1.05% 0.7385 0.18% 0.2915 1.64%
JointMotion 3.2400 -11.75% 1.3830 -9.36% 0.7090 -3.82% 0.2847 -0.73%
HPTR None 2.6003 1.1682 0.6030 0.2331
PreTraM [15] 2.5049 -3.66% 1.0981 -5.99% 0.5863 -2.78% 0.2345 0.60%
JointMotion 2.4006 -7.68% 1.0564 -9.58% 0.5591 -7.28% 0.2297 -1.46%
Wayformer (joint) None 2.3529 1.0209 0.5461 0.2273
JointMotion 2.2823 -3.00% 0.9939 -2.64% 0.5270 -3.50% 0.2143 -5.72%
AV2 HPTR None 2.2550 1.1380 - 0.0988
JointMotion WOMD 2.1530 -4.53% 1.1370 -0.09% - 0.1025 3.75%
@@
Table 2: Comparing scene-level self-supervision methods. All metrics are computed using the Waymo Open Motion interactive (WOMD) and Argoverse 2 Forecasting (AV2) validation splits. Best scores are bold.
Unlike PreTraM, the proposed adaptive pre-training decoder combined with our instance-level objective enable our method to adapt to models with early fusion mechanisms (e.g., Wayformer). Specifically, modality-specific masked polyline modeling with learned queries in the same number as input tokens enables our method to decode modality-specific tokens (i.e., agent, lane, and traffic light) from compressed latent representations. This is particular relevant since the current state-of-the-art methods on the Argoverse 1 Forecasting (ProphNet [11] via AiP tokens) and Argoverse 2 Forecasting (QCNeXt [39] via query-centric modeling) benchmarks rely on fusion mechanisms with compressed latent representations as well.
Furthermore, our pre-training leads to comparable improvements on the interactive and the regular validation splits (cf. Table 1), while pre-training with PreTraM leads to smaller improvements on the interactive validation split. We hypothesise that with our additional instance-level objective, more fine-grained trajectory details are learned, which is more important for close trajectories of interacting agents. The lowest block shows that our method leads to transferable representations. Specifically, pre-training on WOMD improves fine-tuning on AV2.
4.3 Comparison with state-of-the-art methods for joint motion prediction
@@
Experimental setup. We pre-train HPTR (configured as in Section 4.2) for 10 hours using JointMotion and fine-tune for 100 hours. We evaluate the methods in test and validation splits of WOMD. We use the official challenge website to compute performance metrics.
Results. Table 3 presents a comparative analysis of various state-of-the-art methods for joint motion prediction on interactive splits of WOMD. In the test split, we compare four different approaches. The Scene Transformer has a mAP of 0.1192 and is outperformed by the other methods in all metrics. GameFormer achieves a higher mAP of 0.1376 and exhibits competitive performance in terms of minADE and minFDE. MotionDiffuser and JointMotion (HPTR) are close contenders, with MotionDiffuser performing slightly better in all of the metrics compared to JointMotion (HPTR).
For the validation split, the results are similar. GameFormer (joint) scores a mAP of 0.1339, with MotionLM (single replica) outperforming it in terms of mAP, achieving 0.1687. JointMotion (HPTR) presents an improved mAP of 0.1761 compared to its counterparts. Following [21], we compare against the single replica model of MotionLM and show its ensembling version for reference. MotionLM (ensemble) demonstrates superior overall performance, as expected given the increased modeling capacity.
Split Method (config) Venue mAP $\uparrow$ minADE $\downarrow$ minFDE $\downarrow$ MR $\downarrow$ OR $\downarrow$
@@
GameFormer (joint) [40] ICCV’23 0.1376 0.9161 1.9373 0.4531 0.2112
MotionDiffuser [24] CVPR’23 0.1952 0.8642 1.9482 0.4300 0.2004
@@
Val GameFormer (joint) [40] ICCV’23 0.1339 0.9133 1.9251 0.4564 -
@@
Table 3: Comparison with state-of-the-art methods for joint motion prediction. All methods are evaluated on interactive splits of the Waymo Open Motion dataset. Following [21], we compare against single replica versions of joint prediction methods, ensemble versions are shown for reference. Best scores are bold, second best are underlined.